# PC-SRIF: Preconditioned Cholesky-based Square Root Information Filter for Vision-aided Inertial Navigation

### v1

@@ I Introduction
A vision-aided inertial navigation system (VINS), which combines data from inertial measurement units (IMUs) and cameras, has been a primary solution for 6 degrees-of-freedom (DOF) pose (position and orientation) estimation tasks [1]. Given its low cost, independence from external infrastructure such as GPS, and high accuracy, VINS has led to numerous successful applications, including augmented reality (AR), virtual reality (VR), and robotics. In these applications, it is essential to provide accurate pose estimation with minimal latency, typically on a platform with limited computing resources, such as VR headsets and AR glasses. While existing VINS estimators have demonstrated their capability on these platforms, VINS remains an intensive perception task. Thus, it is beneficial to further enhance the efficiency of VINS estimators, so as to conserve battery power and provide low-latency pose estimates with minimal computational resource contention. In what follows, we first provide a concise overview of existing VINS estimators and then focus on those which prioritize efficiency.
@@
IF employs the information matrix rather than the covariance matrix to parameterize the Gaussian distribution, and most optimization-based VINS (e.g., [7, 19, 8, 20, 21]) can be categorized as utilizing IF. For VINS, IF can attain lower complexity if the problem structure is appropriately exploited. Specifically, when $m$ significantly surpasses $n$, IF offers a notable computational advantage over other variants, attributed to employing the Cholesky solver for the linear least-squares (LS) problem arising in the update step [$O(1mn^{2})$ floating point operations (FLOPs)]. However, IF has been criticized for its numerical instability. In particular, [22] asserted that the information matrix in VINS is often ill-conditioned, necessitating double precision (float64) for IF. Thus, the square root information filter (SRIF) is employed instead [22, 23, 24, 25, 26], where the QR decomposition is used for solving the LS, requiring more FLOPs [$O(2mn^{2})$] but sufficiently stable to operate in float32 for VINS.
@@ II Information Filters Review and PC-SRIF
The primary motivation behind introducing preconditioners is to enable the Cholesky LS solver, thereby outperforming the QR LS solver employed by SRIF. Consequently, the computational cost associated with the preconditioners in PC-SRIF must be sufficiently low to not offset the benefits offered by the LS solver. Notably, general preconditioners often lack computational efficiency or effectiveness in improving conditioning. Thus, domain knowledge is essential for designing effective preconditioners. In the subsequent section, we present our PC-SRIF based VINS, where we leverage VINS specific information to address the conditioning issue and ensure numerical stability in float32.
@@ III-C Propagation and State Augmentation
During the propagation step, new states such as the new pose and new features are added to the state vector. The square root information matrix $\mathbf{R}$ is modified accordingly similar to [22]. However, in contrast to [22], which leaves the augmented $\mathbf{R}$ non-upper-triangular and replies on a Householder QR in the following marginalization step, we perform Givens QR to make $\mathbf{R}$ upper-triangular before marginalization. As what the next subsection shows, using Givens QR to keep $\mathbf{R}$ upper-triangular is essential for utilizing sparsity and ensuring efficiency.
@@ III-E Update and Conditioning Analysis
First, note that PC-SRIF only needs to focus on the condition number of $\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}$, a smaller matrix than the $\mathcal{I}^{\oplus}$ or $\mathbf{R}^{\oplus}$ that needs to be considered in IF. This frees us from dealing with other states correlated with the unobservable directions but not involved in the visual measurement model such as ${}^{\scriptscriptstyle G}\mathbf{v}_{\scriptscriptstyle I_{k}}$. In addition, although the goal is to find a matrix $\mathbf{M}$ such that $\kappa(\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}\mathbf{M}^{-1})$ is small, we require $\mathbf{M}$ to be able to be computed quickly, using information already available to the estimator. Thus, we find preconditioners for $\mathbf{R}_{\scriptscriptstyle 22}$ as an approximation of $\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}$. As demonstrated by the experimental findings, this approach is adequate for substantially reducing the condition number in VINS.
@@
$\mathbf{M}_{\scriptscriptstyle Jacobi}(i,j)=\begin{cases}\lVert\mathbf{R}_{\scriptscriptstyle 22s}(:,i)\rVert_{2}&\text{if }i=j\\ 0&\text{else}\end{cases}$ (19)
@@
Note that $\mathbf{M}_{\scriptscriptstyle Jacobi}$ and $\mathbf{M}_{\scriptscriptstyle SPAI}$ are sparse matrices and thus can be constructed and applied without the explicit formation or computing inversion. Thus, our preconditioning introduces minor computational cost compared to the LS solver itself, which we demonstrate in the next section.
@@ IV-A Experiment Setup
The goal of our experiments is to provide a comparative analysis of various filtering-based VINS estimators and validate the theoretical advantages of PC-SRIF on practical platforms with limited computational resources. To this end, we implemented a VINS with three estimator options: KF, SRIF, and PC-SRIF. Except for the estimator, all other modules such as image processing and measurement function evaluation are shared, ensuring a fair comparison among the estimators themselves. In this experiment VINS, the number of poses in the sliding window is set to $l=11$, and at most 15 SLAM features and 35 MSCKF features (as defined in [22]) are processed at each time step. Note that since all these estimators are mathematically equivalent, while they may exhibit different efficiency, the only source of difference in their estimates is numerical error.
We evaluate our VINS on a cell phone dataset collected on a Pixel 3. Additionally, we include the results of a state-of-the-art open source VINS, OpenVINS [11] (a float64 KF), on monocular EuRoC dataset [31]. All systems are executed on a Pixel 7 phone and affiliated to one 2.85 GHz Cortex-X1 CPU. The absolute trajectory error (ATE) is employed for accuracy comparison.
In the subsequent subsections, we first present a numerical stability analysis, which demonstrates the effectiveness of our preconditioners in reducing the condition number and PC-SRIF’s numerical stability. Then, we evaluate the end-to-end accuracy of the PC-SRIF, and the estimator runtime is measured to showcase its superior efficiency.
@@ IV-B Numerical Stability Analysis
The key factor if a Cholesky LS solver can be applied stably is the condition number of the (square root) information matrix given a machine epsilon [32]. As discussed earlier, in solving in LS (7) of SRIF, only a sub-block $\mathbf{R}^{\oplus}_{\scriptscriptstyle 22}$ of the square root information matrix needs to be inverted. Nevertheless, as shown in Fig. 4, with the global pose parameterization, $\kappa^{2}(\mathbf{R}_{\scriptscriptstyle 22}^{\oplus})$ can be as high as $10^{9}$, which necessities float64 for the Cholesky LS solver. $\kappa^{2}(\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}\mathbf{D}^{-1})$ is still high in float32 and increasing with time, suggesting that only applying a diagonal scaling does not resolve the issue. In contrast, in PC-SRIF, the preconditioned square root information matrix $\mathbf{R}_{\scriptscriptstyle 22}\mathbf{M}^{-1}$ and $\mathbf{R}^{\oplus}_{\scriptscriptstyle 22}\mathbf{M}^{-1}$ are constantly well-conditioned, This evidences two conclusions: (i) Our proposed VINS preconditioners are effective for enabling the Cholesky solver in float32. (ii) For constructing VINS preconditioners, it is sufficient to consider only the prior information matrix $\mathbf{R}_{\scriptscriptstyle 22}$ rather than the posterior $\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}$.
@@ IV-C Accuracy and Efficiency
As mentioned before, PC-SRIF differs from KF and SRIF only in numerical difference, and thus we first compare VINS’ accuracy to show PC-SRIF’s numerical stability using float32. As shown in Table I,II, all the estimators implemented by us, regardless in float32 and float64, achieve similar accuracy, which demonstrate that they are all sufficiently stable in float32, and PC-SRIF does not compromise accuracy compared to KF and SRIF. In addition, they are comparable with OpenVINS, indicating that our VINS achieves reasonable accuracy for comparing efficiency.
Next, we evaluate the computational cost of the considered estimators to demonstrate PC-SRIF’s superior efficiency. Table III summarizes the time spent on the estimators, i.e., linear algebra operations in propagation, marginalization, and update, excluding other modules such as image processing (4.65 ms), measurement function/Jacobian evaluation and bookkeeping (1.96 ms). Though KF’s marginalization is trivial, thanks to our novel algorithm, (PC-)SRIF’s marginalization cost is also minor compared to other steps. As expected, the update step is the most expensive for all filters, while PC-SRIF makes it faster by 62% compared to SRIF. In addition, out of the 0.53 ms in the update, the time spent on the preconditioning is 0.08 ms, indicating that its overhead is minor. Overall, KF is much slower than SRIF due to its expensive update, agreeing with [22]’s results. Remarkably, PC-SRIF is even faster than SRIF by 41%, which validates the theoretical advantages in our previous analysis.
TABLE I: EuRoC dataset (all 11 sequences): Mean ATE in meters/degrees
@@
TABLE II: Cell phone dataset: Mean ATE in meters/degrees

### v2

@@ I Introduction
A vision-aided inertial navigation system (VINS), which combines data from inertial measurement units (IMU) and cameras, has been a primary solution for 6 degrees-of-freedom (DOF) pose (position and orientation) estimation tasks [1]. Given its low cost, independence from external infrastructure such as GPS, and high accuracy, VINS has led to numerous successful applications, including augmented reality (AR), virtual reality (VR), and robotics. In these applications, it is essential to provide accurate pose estimation with minimal latency, typically on a platform with limited computing resources, such as VR headsets and AR glasses. While existing VINS estimators have demonstrated their capability on these platforms, VINS remains an intensive perception task. Thus, it is beneficial to further enhance the efficiency of VINS estimators, so as to conserve battery power and provide low-latency pose estimates with minimal computational resource contention. In what follows, we first provide a concise overview of existing VINS estimators and then focus on those which prioritize efficiency.
@@
IF employs the information matrix rather than the covariance matrix to parameterize the Gaussian distribution, and most optimization-based VINS (e.g., [7, 19, 20, 21]) can be categorized as utilizing IF, in the sense that their prior information from marginalization includes an information matrix. For VINS, IF can attain lower complexity if the problem structure is appropriately exploited. Specifically, when $m$ significantly surpasses $n$, IF offers a notable computational advantage over other variants, attributed to employing the Cholesky solver for the linear least-squares (LS) problem arising in the update step [$O(1mn^{2})$ floating point operations (FLOPs)]. However, IF has been criticized for its numerical instability. In particular, [22] asserted that the information matrix in VINS is often ill-conditioned, necessitating double precision (float64) for IF. Thus, the square root information filter (SRIF) is employed instead [22, 23, 24, 25, 26], where the QR decomposition is used for solving the LS, requiring more FLOPs [$O(2mn^{2})$] but sufficiently stable to operate in float32 for VINS.
@@ II Information Filters Review and PC-SRIF
The primary motivation behind introducing preconditioners is to enable the Cholesky LS solver, thereby outperforming the QR LS solver employed by SRIF. Consequently, the computational cost associated with the preconditioners in PC-SRIF must be sufficiently low to not offset the benefits offered by the LS solver. Notably, general preconditioners often lack computational efficiency or effectiveness in improving conditioning, compared to those employing domain knowledge. In the subsequent section, we present our PC-SRIF based VINS, where we leverage VINS specific information to design effective preconditioners and ensure numerical stability in float32.
@@ III-C Propagation and State Augmentation
During the propagation step, new states such as the new pose and new features are added to the state vector. The square root information matrix $\mathbf{R}$ is modified accordingly similar to [22]. However, in contrast to [22], which leaves the augmented $\mathbf{R}$ non-upper-triangular and relies on a Householder QR in the following marginalization step, we perform Givens QR to make $\mathbf{R}$ upper-triangular before marginalization. As what the next subsection shows, using Givens QR to keep $\mathbf{R}$ upper-triangular is essential for utilizing sparsity and ensuring efficiency.
@@ III-E Update and Conditioning Analysis
First, note that PC-SRIF only needs to focus on the condition number of $\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}$, a smaller matrix than the $\mathcal{I}^{\oplus}$ or $\mathbf{R}^{\oplus}$ that needs to be considered in IF. This frees us from dealing with other states correlated with the unobservable directions but not involved in the visual measurement model such as ${}^{\scriptscriptstyle G}\mathbf{v}_{\scriptscriptstyle I_{k}}$. Second, although the goal is to find a matrix $\mathbf{M}$ such that $\kappa(\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}\mathbf{M}^{-1})$ is small, to limit the cost, we use information already available before constructing $\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}$. Thus, we find preconditioners for $\mathbf{R}_{\scriptscriptstyle 22}$ as an approximation. As demonstrated by the experiments, this approach is adequate for substantially reducing the condition number in VINS.
@@
$\mathbf{M}_{\scriptscriptstyle Jacobi}(i,j)=\lVert\mathbf{R}_{\scriptscriptstyle 22s}(:,i)\rVert_{2}\text{ if }i=j\text{ else }0$ (19)
@@
Note that $\mathbf{M}_{\scriptscriptstyle Jacobi}$ and $\mathbf{M}_{\scriptscriptstyle SPAI}$ are matrices with known sparse patterns, and thus can be applied without explicit construction or computing inversion. Consequently, our preconditioning introduces minor computational cost to the LS solver, which we demonstrate in the next section.
@@ IV-A Experiment Setup
The goal of our experiments is to provide a comparative analysis of various filtering-based VINS estimators and validate the theoretical advantages of PC-SRIF on practical platforms with limited computational resources. To this end, we implemented a VINS with three estimator options: KF, SRIF, and PC-SRIF. Except for the estimator, all other modules such as image processing and measurement function evaluation are shared, ensuring a fair comparison among the estimators themselves. In this VINS, the number of poses in the sliding window is set to $l=11$, and at most 15 SLAM features and 35 MSCKF features (as defined in [22]) are processed at each time step. Note that since all these estimators are mathematically equivalent, while they may exhibit different efficiency, the only source of difference in their estimates is numerical error.
We evaluate our VINS on a cell phone dataset collected on a Pixel 3. Additionally, we include the results of a state-of-the-art open source VINS, OpenVINS [11] (a float64 KF), on monocular EuRoC dataset [31]. All systems are executed on a Pixel 7 phone and affiliated to one 2.85 GHz Cortex-X1 CPU. The absolute trajectory error (ATE) and relative trajectory error (RTE, with 1 sec interval) are employed for accuracy comparison.
The following subsections will first present a numerical stability analysis, demonstrating how our preconditioners ensure PC-SRIF’s numerical stability by reducing the condition number. Subsequently, we will evaluate the end-to-end accuracy of the PC-SRIF to confirm that its efficiency gains do not compromise accuracy. Finally, the estimator runtime will be measured to showcase its superior efficiency.
@@ IV-B Numerical Stability Analysis
The key factor if a Cholesky LS solver can be applied stably is the condition number of the (square root) information matrix given a machine epsilon [32]. As discussed earlier, in solving in LS (7) of SRIF, only a sub-block $\mathbf{R}^{\oplus}_{\scriptscriptstyle 22}$ of the square root information matrix needs to be inverted. Nevertheless, as shown in Fig. 4, with the global pose parameterization, $\kappa^{2}(\mathbf{R}_{\scriptscriptstyle 22}^{\oplus})$ can be as high as $10^{9}$, which necessitates float64 for the Cholesky LS solver. $\kappa^{2}(\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}\mathbf{D}^{-1})$ is still high in float32 and increases with time, suggesting that only applying a diagonal scaling does not resolve the issue. In contrast, in PC-SRIF, the preconditioned square root information matrix $\mathbf{R}_{\scriptscriptstyle 22}\mathbf{M}^{-1}$ and $\mathbf{R}^{\oplus}_{\scriptscriptstyle 22}\mathbf{M}^{-1}$ are constantly well-conditioned, This evidences two conclusions: (i) Our proposed VINS preconditioners are effective for enabling the Cholesky solver in float32. (ii) For constructing VINS preconditioners, it is sufficient to consider only the prior information matrix $\mathbf{R}_{\scriptscriptstyle 22}$ rather than the posterior $\mathbf{R}_{\scriptscriptstyle 22}^{\oplus}$.
@@ IV-C Accuracy and Efficiency
As mentioned before, PC-SRIF differs from KF and SRIF only in numerical difference, and thus we first compare VINS’ accuracy to show PC-SRIF’s numerical stability in float32. As shown in Table I,II, all the estimators implemented by us, regardless in float32 and float64, achieve similar accuracy, demonstrating that they are all sufficiently stable in float32. PC-SRIF does not compromise accuracy compared to SRIF even if it employs the less stable Cholesky solver, attributed to its improved conditioning. In addition, they are comparable with OpenVINS, indicating that our VINS achieves reasonable accuracy for comparing efficiency.
Next, we evaluate the computational cost of the considered estimators to demonstrate PC-SRIF’s superior efficiency. Table III summarizes the time spent on the estimators, i.e., linear algebra operations in propagation, marginalization, and update, excluding other modules such as image processing, measurement function evaluation, and bookkeeping. Though KF’s marginalization is trivial, thanks to our novel algorithm, (PC-)SRIF’s marginalization cost is minor compared to other steps. As expected, the update step is the most expensive for all, while PC-SRIF makes it faster by 62% compared to SRIF. In addition, out of the 0.53 ms in the update, the time spent on the preconditioning is only 0.08 ms, a minor overhead. Overall, KF is much slower than SRIF due to its expensive update, agreeing with [22]’s results. Remarkably, PC-SRIF is even faster than SRIF by 41%, which validates the theoretical advantages in our previous analysis.
TABLE I: EuRoC dataset (11 sequences): Mean ATE (top 2 rows) and RTE (bottom 2 rows) in meters/degrees
@@
float32 0.046 / 0.33 0.045 / 0.31 0.045 / 0.31 -
float64 0.047 / 0.33 0.046 / 0.31 0.045 / 0.31 0.042 / 0.30
TABLE II: Cell phone dataset: Mean ATE (top 2 rows) and RTE (bottom 2 rows) in meters/degrees